# Catalyst: Out-of-Distribution Detection via Elastic Scaling

### v1

@@ 4.6 Comparison with Other Baselines
Comparing Catalyst against three contemporary methods, AdaScale [50], NCI [35] and fDBD [34], confirms its superiority, particularly when used as a synergistic module. Aginst, AdaScale’s reported reslults using DenseNet-101 on CIFAR-100, ${\texttt{Catalyst}}(m)+\texttt{ReAct}$ outperforms AdaScale yielding a 32.45% gain over the best AdaScale variant. Against NCI’s reported results (which were obtained on the OpenOOD settings), ${\texttt{Catalyst}}(m)+\texttt{ReAct}$ achieves the average FPR95 by a significant 33.43% on CIFAR-10 (ResNet-18). This advantage is even more pronounced against fDBD, where our method achieves a substantial FPR95 reduction of 65.54% on ImageNet (ResNet-50). We also provide a detailed comparison with 19 existing OOD detection methods in literature in Appendix F.
@@ 5.1 Choice of Layer for Computing $\gamma$
This finding is intuitively aligned with the principles of hierarchical feature learning [47, 71]. These initial layers learn general, low-level features such as edges, textures, and color blobs, which are fundamental properties shared by all natural images. Since both ID and OOD samples contain these common features, their activation statistics in these early layers are highly similar, resulting in the non-discriminative, overlapping $\gamma$ distributions we observed. In sharp contrast, the distribution from the final residual stage (Layer 4), immediately preceding GAP, provides a better separation, because it is trained to recognize the complex, high-level concepts and structures specific to the ID classes, which OOD samples lacks. This analysis, which held true across all tested OOD datasets and architectures, empirically validates our focus: the penultimate layer’s pre-pooling feature map is not a layer of convenience but the most reliable source of a potent signal for Catalyst. The complete details are presented in Appendix H.
@@ 5.4 Scaling Factor $\gamma$ as a Scoring Metric
We conducted an analysis to determine if $\gamma$ is powerful enough to serve as a standalone OOD score, similar to MSP [16] or Energy [36]. Our findings show that $\gamma$ computed from std ($\gamma_{\text{std}}$) and max ($\gamma_{\text{max}}$) are consistently robust signals. On both the CIFAR benchmarks (Table 21) and the large-scale ImageNet benchmark (Table 20), these two statistics is consistently better than Energy baseline, proving they are viable and generalizable standalone scores. In contrast, the entropy provides a critical insight. While $\gamma_{\text{entropy}}$ appears to be the distinguishable signal on CIFAR, this trend is inconsistent on ImageNet. On this more complex benchmark, $\gamma_{\text{entropy}}$ fails to generalize, suffering a performance collapse and lagging behind Energy. This analysis confirms that entropy, while potent in some cases is not a reliable or generalizable statistic for a robust, all-purpose method. The complete details are provided in Appendix K.
6 Limitations and Future Work
@@
Additionally, we limit the scope of Catalyst to CNN-based architectures. The reason behind this is two-fold: (a) Competitive baselines in the literature [36, 33, 44, 62, 54, 55, 5, 67] extensively use CNN-based architectures. For fair comparison, we adopt similar architectures to evaluate Catalyst. (b) CNN-based architectures continue to be widely used in both the research community and real-world applications. A comprehensive benchmark study carried out in prior work [12] has shown convolutional networks such as ResNet [13] and ConvNeXt [37, 65] remain the default choice in real-world vision systems (including object detection, segmentation, retrieval, and classification) due to their strong inductive bias (translation invariance), computational efficiency, strong performance on moderate-scale data, and extensive ecosystem of pretrained models.
The core principle of our method – leveraging statistical cues from penultimate pre-pooled activation map – is a general strategy that can be extended beyond CNNs to architectures like Vision Transformers (ViTs) [6]. However, adapting Catalyst to derive an effective scaling factor $\gamma$ from the intermediate blocks of a transformer requires substantial research and engineering. We identify the extension of our framework to transformer-based models as a promising and significant direction for future work.
@@ Appendix A Description of Baseline Methods
    $h^{\text{SCALE}}(\mathbf{x})=\exp(r)\cdot h(\mathbf{x}).$
@@ E.1 Near-OOD Evaluation
We also evaluate Catalyst on the challenging near-OOD task of distinguishing CIFAR-10 from CIFAR-100, a commonly used setup used in prior work [11]. As shown in Table 8, while the separation is inherently more difficult for all methods, Catalyst still provides a performance improvement over the baselines when applied in tandem, demonstrating its robustness even in fine-grained detection scenarios. For instance, with ResNet-18, ${\texttt{Catalyst}}(m)+\texttt{ReAct}$ reduces the FPR95 from 52.04% to 49.62%, an improvement of 4.65%.
@@ G.1 Code and Data Availability
The complete source code for used in Catalyst, along with the scripts used to run all experiments and generate figures, is publicly available on GitHub. [^2] https://github.com/epsilon-2007/Catalyst We will also provide the model weights for our trained CIFAR models. All datasets used in this work (CIFAR-10, CIFAR-100, ImageNet-1k, and all OOD benchmarks) are publicly available and were used without modification, following the standard preprocessing steps described in their original publications and common benchmarks.
@@ G.2 Hyperparameter Selection
The clipping threshold $\boldsymbol{c}$ (Eq. 3) is crucial for enhanced performance, as it must be set to optimally distinguish ID from OOD data. Analogous to ReAct [54], we do not tune $\boldsymbol{c}$ directly; instead, we control it by setting it to the $p$-th percentile of the ID activation distribution (e.g, when $p=95$, it indicates that 95% percent of the ID activations are less than the threshold $\boldsymbol{c}$). The choice of this percentile $p$ is the key hyperparameter to be tuned. To select the optimal $p$, we follow established protocols from prior work [55, 54] and create a proxy OOD validation set, which is generated by adding pixel-wise Gaussian noise $\mathcal{N}(0,0.2)$ to images from the ID validation set. We then select the percentile $p$ that yields the best OOD separation on this proxy task. This two-step procedure – using a percentile for the mechanism and a proxy set for tuning – is a robust tuning strategy grounded in prior work. The selected $p$ values are:
@@
  • ImageNet: For our default method (${\texttt{Catalyst}}+\text{Energy}$), a single percentile of $p=75$ is used for all ImageNet models. When combining with baseline like ReAct (${\texttt{Catalyst}}+\texttt{ReAct}$), we found it optimal to use a single, shared percentile $p$ across all three statistics (mean, std, and max). The optimal shared percentile $p$ varies by model: $p=15$ for ResNet-34 and ResNet-50, $p=35$ for MobileNet-v2, and $p=52$ for DenseNet-121.
As our hyperparameter search for ImageNet demonstrated, the optimal shared percentile $p$ varies across different architectures (e.g., $p=15$ for ResNet-50 vs. $p=52$ for DenseNet-121). This empirical finding is highly intuitive and aligns with our method’s design. The optimal $p$ (which sets the clipping threshold $c$) is naturally coupled with the model’s architecture, particularly the dimension ($n$) of the penultimate layer. This is because our scaling factor $\gamma$ (Equation 4) is an aggregation (a sum) over all $n$ channels. A model with a larger channel dimension (e.g., ResNet-50, $n=2048$) will produce a sum of a very different magnitude than a model with a smaller dimension (e.g., DenseNet-121, $n=1024$). Therefore, a different clipping percentile $p$ is required for each architecture and produce the most discriminative $\gamma$ signal.
@@ G.3 Baseline Hyperparameter Tuning
When re-evaluating these baselines on new architectures not present in their original work, we performed a new hyperparameter search using the same validation procedures and search spaces they described. This ensures that every baseline is as strong as possible for each specific model. Key hyper-parameters for these methods are summarized below:
@@ Appendix I Analysis of Fusion Strategy
$S^{*}(\mathbf{x};\theta,\gamma)$ $=\gamma(\mathbf{x};f)+S(\mathbf{x};\theta)$ (12b)
@@
$\gamma^{*}(\mathbf{x};f)=\sum_{i=1}^{n}\min(f_{i}(\mathbf{x}),c)$ (13a)
@@
For proposed multiplicative strategy, the optimal threshold $c^{*}$ is set by selecting a moderate percentile as detailed in reproducibility section in Appendix G of the ID training data’s $f_{i}(\mathbf{x})$ values. This procedure is robust, stable, and aligns with foundational post-hoc methods like ReAct and SCALE (details in Appendix G). For the additive strategy, we empirically found that the optimal threshold $c^{+}$ was consistently an order of magnitude smaller, (e.g., $c^{+}\approx 0.1\times c^{*}$). On ImageNet, this optimal $c^{+}$ value corresponds to an extremely low percentile of the ID data (e.g., less than 1st for ResNet-50).
@@
Computing the Scaling Factor $\gamma$. The $\gamma(\mathbf{x})$ computation using the median follows the same principle described in Section 3: a higher median activation is assumed to indicate an ID sample. The Shannon entropy, however, exhibits the opposite behavior. As alluded to in our motivation (Section 1) and empirically demonstrated (Figure 1), ID samples typically have lower entropy (i.e., less uncertainty) than OOD samples. So, to maintain the convention that a ID sample exhibits higher score than OOD samples, the entropy-based scaling factor must be inverted (i.e, $\frac{1}{\gamma(\mathbf{x})}$ ) before it is applied to the baseline score.
@@
The analysis of Shannon entropy is more nuanced and reveals a critical insight. Entropy is not consistently ineffective; rather, it is inconsistent. In specific cases, entropy can be very effective. For example, on the ImageNet benchmark (Table 23), the entropy-based $\gamma$ improves the ReAct+DICE baseline by 14.65% on the MobileNet-V2 architecture, achieving the best performance for that specific model. However, this strong performance is not generalizable. On the same dataset but with a ResNet-50 backbone (Table 22), the improvement is minimal (5.90%) and lags behind our proposed mean/std/max combination. This inconsistency aligns can also be seen in ablation study of using scaling factor standalone as scoring metric in Appendix K, where we show that $\gamma_{\text{entropy}}$ as a standalone score was dominant on CIFAR but failed to generalize to ImageNet.

### v2

@@ 4.6 Comparison with Other Baselines
Comparing Catalyst against three contemporary methods, AdaScale [50], NCI [35] and fDBD [34], confirms its superiority, particularly when used as a complementary module. Aginst, AdaScale’s reported reslults using DenseNet-101 on CIFAR-100, ${\texttt{Catalyst}}(m)+\texttt{ReAct}$ outperforms AdaScale yielding a 32.45% gain over the best AdaScale variant. Against NCI’s reported results (which were obtained on the OpenOOD settings), ${\texttt{Catalyst}}(m)+\texttt{ReAct}$ achieves the average FPR95 by a significant 33.43% on CIFAR-10 (ResNet-18). This advantage is even more pronounced against fDBD, where our method achieves a substantial FPR95 reduction of 65.54% on ImageNet (ResNet-50). We also provide a detailed comparison with 19 existing OOD detection methods in literature in Appendix F.
@@ 5.1 Choice of Layer for Computing $\gamma$
This finding is intuitively aligned with the principles of hierarchical feature learning [47, 71]. These initial layers learn general, low-level features such as edges, textures, and color blobs, which are fundamental properties shared by all natural images. Since both ID and OOD samples contain these common features, their activation statistics in these early layers are highly similar, resulting in the non-discriminative, overlapping $\gamma$ distributions we observed. In sharp contrast, the distribution from the final residual stage (Layer 4), immediately preceding GAP, provides a better separation, because it is trained to recognize the complex, high-level concepts and structures specific to the ID classes, which OOD samples lack. This analysis, which held true across all tested OOD datasets and architectures, empirically validates our focus: the penultimate layer’s pre-pooling feature map is not a layer of convenience but the most reliable source of a potent signal for Catalyst. The complete details are presented in Appendix H.
@@ 5.4 Scaling Factor $\gamma$ as a Scoring Metric
We conducted an analysis to determine if $\gamma$ is powerful enough to serve as a standalone OOD score, similar to MSP [16] or Energy [36]. Our findings show that $\gamma$ computed from std ($\gamma_{\text{std}}$) and max ($\gamma_{\text{max}}$) are consistently robust signals. On both the CIFAR benchmarks (Table 21) and the large-scale ImageNet benchmark (Table 20), these two statistics are consistently better than Energy baseline, proving they are viable and generalizable standalone scores. In contrast, the entropy provides a critical insight. While $\gamma_{\text{entropy}}$ appears to be the distinguishable signal on CIFAR, this trend is inconsistent on ImageNet. On this more complex benchmark, $\gamma_{\text{entropy}}$ fails to generalize, suffering a performance collapse and lagging behind Energy. This analysis confirms that entropy, while potent in some cases is not a reliable or generalizable statistic for a robust, all-purpose method. The complete details are provided in Appendix K.
6 Scope and Future Work
@@
Additionally, we provide a comprehensive analysis focused on CNN-based architectures, with Catalyst applied within this setting. This focus is motivated by two factors: (a) Competitive baselines in the literature [36, 33, 44, 62, 54, 55, 5, 67] extensively use CNN-based architectures. For fair comparison, we adopt similar architectures to evaluate Catalyst. (b) CNN-based architectures continue to be widely used in both the research community and real-world applications. A comprehensive benchmark study carried out in prior work [12] has shown convolutional networks such as ResNet [13] and ConvNeXt [37, 65] remain the default choice in real-world vision systems (including object detection, segmentation, retrieval, and classification) due to their strong inductive bias (translation invariance), computational efficiency, strong performance on moderate-scale data, and extensive ecosystem of pretrained models.
The core principle of our method, leveraging statistical cues from penultimate pre-pooled activation map is a general strategy that can be extended beyond CNNs to architectures like Vision Transformers (ViTs) [6]. However, adapting Catalyst to derive an effective scaling factor $\gamma$ from the intermediate blocks of a transformer requires substantial research and engineering. We are actively exploring the extension of our framework to transformer-based models.
@@ Appendix A Description of Baseline Methods
    $h^{\text{SCALE}}(\mathbf{x};\theta)=\exp(r)\cdot h(\mathbf{x};\theta).$
@@ E.1 Near-OOD Evaluation
We also evaluate Catalyst on the challenging near-OOD task of distinguishing CIFAR-10 from CIFAR-100, a commonly used setup used in prior work [11]. As shown in Table 8, while the separation is inherently more difficult for all methods, ${\texttt{Catalyst}}(m)$ still provides a marginal performance improvement over the baselines when applied in tandem, demonstrating its robustness even in fine-grained detection scenarios. For instance, with ResNet-18, ${\texttt{Catalyst}}(m)+\texttt{ReAct}$ reduces the FPR95 from 52.04% to 49.62%, an improvement of 4.65%. We attribute the limited improvement in near-OOD settings to the high similarity of the learned penultimate representations. A valuable direction for future research is to design a suitable scaling factor $\gamma$ in near-ood evaluation settings.
@@ G.1 Code and Data Availability
The complete source code used in Catalyst, along with the scripts used to run all experiments and generate figures, is publicly available on GitHub. [^2] https://github.com/bingabid/Catalyst We will also provide the model weights for our trained CIFAR models. All datasets used in this work (CIFAR-10, CIFAR-100, ImageNet-1k, and all OOD benchmarks) are publicly available and were used without modification, following the standard preprocessing steps described in their original publications and common benchmarks.
@@ G.2 Hyperparameter Selection
The clipping threshold $\boldsymbol{c}$ (Eq. 3) is crucial for enhanced performance, as it must be set to optimally distinguish ID from OOD data. Analogous to ReAct [54], we do not tune $\boldsymbol{c}$ directly; instead, we control it by setting it to the $p$-th percentile of the ID activation distribution (e.g, when $p=95$, it indicates that 95% of the ID activations are less than the threshold $\boldsymbol{c}$). The choice of this percentile $p$ is the key hyperparameter to be tuned. To select the optimal $p$, we follow established protocols from prior work [55, 54] and create a proxy OOD validation set, which is generated by adding pixel-wise Gaussian noise $\mathcal{N}(0,0.2)$ to images from the ID validation set. We then select the percentile $p$ that yields the best OOD separation on this proxy task. This two-step procedure – using a percentile for the mechanism and a proxy set for tuning – is a robust tuning strategy grounded in prior work. The selected $p$ values are:
@@
  • ImageNet: For our default method (${\texttt{Catalyst}}+\text{Energy}$), a single percentile of $p=75$ is used for all ImageNet models. When combining with baselines like ReAct (${\texttt{Catalyst}}+\texttt{ReAct}$), we found it optimal to use a single, shared percentile $p$ across all three statistics (mean, std, and max). The optimal shared percentile $p$ varies by model: $p=15$ for ResNet-34 and ResNet-50, $p=35$ for MobileNet-v2, and $p=52$ for DenseNet-121.
As our hyperparameter search for ImageNet demonstrated, the optimal shared percentile $p$ varies across different architectures (e.g., $p=15$ for ResNet-50 vs. $p=52$ for DenseNet-121). This empirical finding is highly intuitive and aligns with our method’s design. The optimal $p$ (which sets the clipping threshold $c$) is naturally coupled with the model’s architecture, particularly the dimension ($n$) of the penultimate layer. This is because our scaling factor $\gamma$ (Equation 4) is an aggregation (a sum) over all $n$ channels. A model with a larger channel dimension (e.g., ResNet-50, $n=2048$) will produce a sum of a very different magnitude than a model with a smaller dimension (e.g., DenseNet-121, $n=1024$). Therefore, a different clipping percentile $p$ is required for each architecture to produce the most discriminative $\gamma$ signal.
@@ G.3 Baseline Hyperparameter Tuning
When re-evaluating these baselines on new architectures not present in their original work, we performed a new hyperparameter search using the same validation procedures and search spaces they described. This ensures that every baseline is as strong as possible for each specific model. Key hyperparameters for these methods are summarized below:
@@ Appendix I Analysis of Fusion Strategy
$S^{+}(\mathbf{x};\theta,\gamma)$ $=\gamma(\mathbf{x};f)+S(\mathbf{x};\theta)$ (12b)
@@
$\gamma^{*}(\mathbf{x};f)=\sum_{i=1}^{n}\min(f_{i}(\mathbf{x}),c^{*})$ (13a)
@@
For proposed multiplicative strategy, the optimal threshold $c^{*}$ is set by selecting a moderate percentile as detailed in reproducibility section in Appendix G of the ID training data’s $f_{i}(\mathbf{x})$ values. This procedure is robust, stable, and aligns with foundational post-hoc methods like ReAct and SCALE (details in Appendix G). For the additive strategy, we empirically found that the optimal threshold $c^{+}$ was consistently an order of magnitude smaller, (e.g., $c^{+}\approx 0.1\times c^{*}$). On ImageNet, this optimal $c^{+}$ value corresponds to an extremely low percentile of the ID data (e.g., below the 1st percentile for ResNet-50).
@@
Computing the Scaling Factor $\gamma$. The $\gamma(\mathbf{x})$ computation using the median follows the same principle described in Section 3: a higher median activation is assumed to indicate an ID sample. The Shannon entropy, however, exhibits the opposite behavior. As alluded to in our motivation (Section 1) and empirically demonstrated (Figure 1), ID samples typically have lower entropy (i.e., less uncertainty) than OOD samples. So, to maintain the convention that an ID sample exhibits higher score than OOD samples, the entropy-based scaling factor must be inverted (i.e, $\frac{1}{\gamma(\mathbf{x})}$ ) before it is applied to the baseline score.
@@
The analysis of Shannon entropy is more nuanced and reveals a critical insight. Entropy is not consistently ineffective; rather, it is inconsistent. In specific cases, entropy can be very effective. For example, on the ImageNet benchmark (Table 23), the entropy-based $\gamma$ improves the ReAct+DICE baseline by 14.65% on the MobileNet-V2 architecture, achieving the best performance for that specific model. However, this strong performance is not generalizable. On the same dataset but with a ResNet-50 backbone (Table 22), the improvement is minimal (5.90%) and lags behind our proposed mean/std/max combination. This inconsistency can also be seen in ablation study of using scaling factor standalone as scoring metric in Appendix K, where we show that $\gamma_{\text{entropy}}$ as a standalone score was dominant on CIFAR but failed to generalize to ImageNet.